Calculate the volume (in mL) of a 0.370 M KOH solution that should be added to 4.000 g of HEPES (MW=238.306 g/mol, pKa = 7.56) to give a pH of 7.32.
Answer #1
Concepts and reason

The Henderson-Hassebalch equation is used to determine the pH of a buffer and the concentration of an acid and base.

Fundamentals

Using this equation, the initial volume of the acid or base can be found with the given information

The Henderson-Hassebalch equation is as follows:

pH = pK, +log 3
where,
TA] = molar concentration of conjugate base
[HA] = molar concentration of a undissociated weak acid
PK

Related formulae:

mass
number of moles =
molar mass
numer of moles = concentration volume

molarity of KOH = 0.370M
mass of HEPES = 4.0g
molecular weight of HEPES = 238.306g/mol
pKa of HEPES = 7.56
pH = 7.32

Let HEPES = HA (acid)
hence, A is the conjugate base
mass of HA
initial moles of HA = .
molar mass of HA
4.0g
238.306g/mol
=

moles of KOH added = volume of KOHin ml
1000mL/L
x molarity of KOH
amLx0.370M
1000mL/L
= 0.00037a mol

reaction between HEPES and KOH is as follows:
HA + KOH K + A+H2O
moles of HA left = initial moles of HA - moles of KOH added

pH = pka + log
(moles of A)
pH = pka + log
(moles of HA)
(0.00037a)
7.32= 7.56+log T0 016785-0.00037a)

108 10.016
= -0.24
(0.00037a)
(0.016785-0.00037a)
(0.00037a)
(0.016785-0.00037a)
(0.00037a)
= 0.57544
(0.016785-0.00037a)
= 1

0.00037a = 0.00965876-0.0002129a
0.0005829a = 0.00965876
a = 16.57

Ans:

The volume of KOH in mL = 16.57mL